A cart loaded with sand moves along horizontal road under action of a constant force 5N.In the process sand spills through a hole in the bottom at rate R=0.5kg/s.Find acceleration of the cart at time t=10s after starting. Assume at t=0 mass of cart including sand is 10kg and its velocity is zero.

Grade:12th pass

1 Answers

Shobhit Varshney IIT Roorkee
9 years ago
Dear student,

F = d (mv) / dt = (dm/dt) v + m (dv/dt)
F, dm/dt , v at t=10 s is given.

Calculate dv/dt from this equation.

Good luck in your studies and study hard and achieve heights of success.
Thanks.